OpenStudy (anonymous):

Find the integral of (2 du)/(1+3u). (Answer: (2/3)*ln abs(1+3u)+C where abs is absolute value.)

OpenStudy (anonymous):

\[\int\frac{2}{1+3u}~du\] Substitute \(t=1+3u\), then \(dt=3~du\), or \(\dfrac{1}{3}dt=du\): \[\frac{2}{3}\int\frac{dt}{t}\]

OpenStudy (anonymous):

And then?

OpenStudy (anonymous):

I get it now, thanks.

OpenStudy (anonymous):

You're welcome, sorry about the delay.